OpenStudy (anonymous):

help please i need step by step shown: the line passing through (2,b) and (4,6) is parallel to the line joining (1,4) and (-1,7) find b.

OpenStudy (anonymous):

@tomo

OpenStudy (anonymous):

find the slope of each line. set the two sides equal to each other to find b. slope = (y2-y1)/(x2-x1) using (2,b) and (4,6) (x1,y1) = (2,b) (x2,y2) = (4,6) (6-b)/(4-2) (6-b)/2. that is the slope of the line with (2,b) and (4,6) as two points on the line. find the slope of the other line, then post what you get. then we can go from there.

OpenStudy (anonymous):

the other is 7-4/-1-1= 3/-2 so what do i need to do from here?

OpenStudy (anonymous):

ok, we know that the two lines are parallel, therefore their slopes are equal. set their slopes equal and solve for b. (6-b)/2 = (3/-2)

OpenStudy (anonymous):

what would the 6-b/2 be?

OpenStudy (anonymous):

you want to cross multiply. (6-b)/2 = (3/-2) (6-b)*(-2) = (2*3) -12+2b = 6 2b = 18 b = 9

OpenStudy (anonymous):

oh okay i thought i had to multiply as is im sorry. Thanks.

OpenStudy (anonymous):

yw